1000 identical drops of mercury are charged to a potential 1V each.They join to form a single drop.Find the potenital of this drop? 
 
 

Grade:Select Grade

1 Answers

Kunal
9 years ago
Let r be the radius of small drops and big drop R.
kq/r = 1  ….......(1)
1000* 4/3pi r^3 = 4/3piR^3  ….......(2)
 
potential of big drop = kQ/R                    
 
Substitute R in terms of r using eq 2 & find the ans